OpenStudy (anonymous):

What is the measure of side AB? 10, 9, 98, or, none of these?

OpenStudy (anonymous):



OpenStudy (anonymous):

do you know what kind of triangle that is?

OpenStudy (anonymous):

isosceles right?

OpenStudy (anonymous):

actually it's a right triangle because of the little symbol in the bottom, the 90 degree angle

OpenStudy (anonymous):

yeah.

OpenStudy (mathmath333):

do u know pythgorus formula

OpenStudy (anonymous):

yes a2+b2=c2

OpenStudy (mathmath333):

so what is a and b here

OpenStudy (anonymous):

98

OpenStudy (mathmath333):

no

OpenStudy (mathmath333):

\(\large\tt \color{black}{c^2=98}\)

OpenStudy (anonymous):

oh, well then I guess it would be 49

OpenStudy (mathmath333):

how?

OpenStudy (anonymous):

because I took the sides of both, which are 7 and multiplied it by the other side, which is 7 and got 49

OpenStudy (mathmath333):

In a right angled triangle: the square of the hypotenuse is equal to the sum of the squares of the other two sides

OpenStudy (mathmath333):

\(\large\tt \color{black}{c^2=c\times c}\)

OpenStudy (mathmath333):

\(\large\tt \color{black}{c^3=c\times c\times c}\)

OpenStudy (anonymous):

so 14

OpenStudy (mathmath333):

\(\large\tt \color{black}{a^2+b^2=c^2}\)

OpenStudy (mathmath333):